# Geometric calculations on probability manifolds from reciprocal relations in Master equations

### v3

@@ 3.1. Levi-Civita connection, parallel transport and geodesics
Given $\Phi\in\mathbb{R}^{n}$, we define a vector field $\mathsf{V}_{\Phi}=L(\theta)\Phi\in{T}_{p}\mathcal{P}_{+}$. Suppose $F\in C^{\infty}(\mathcal{P}_{+};\mathbb{R})$, denote the directional derivative of $F$ at direction $\mathsf{V}_{\Phi}$ as
@@
$\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}]:=L(\mathsf{V}_{\Phi_{1}}\theta)\Phi_{2}-L(V_{\Phi_{2}}\theta)\Phi_{1}.$ (13)
@@
where the last equality follows that $\nabla^{2}_{pp}F$ is a symmetric matrix. In detail,
@@
$\bar{\nabla}_{\mathsf{V}_{\Phi_{1}}}\mathsf{V}_{\Phi_{2}}:=\frac{1}{2}\Big(L(\mathsf{V}_{\Phi_{1}}\theta)\Phi_{2}-L(\mathsf{V}_{\Phi_{2}}\theta)\Phi_{1}+L(\theta)\Gamma(\Phi_{1},\Phi_{2},p)\Big).$ (14)
@@
We next compute the parallel transport. Let $\gamma\colon[0,T]\rightarrow\mathcal{P}_{+}$ be a smooth curve, with a parameter $t>0$. Denote $\mathsf{V}_{\Phi}$ as the tangent direction of the curve $\gamma(t)$ at time $t$. I.e., $\frac{d\gamma(t)}{dt}=L(\theta(\gamma(t)))\Phi(t)=\mathsf{V}_{\Phi(t)}$. Consider a vector field $\mathsf{V}_{\eta}$ given by $\eta(t)\in\mathbb{R}^{n}$. Then equation for $\mathsf{V}_{\eta}$ to be parallel along $\gamma(t)$ satisfies
@@ 3.2. Riemannian curvature tensor
$(abcd)=\Phi_{a}^{\mathsf{T}}L(\mathsf{V}_{\Phi_{b}}\mathsf{V}_{\Phi_{c}}\theta)\Phi_{d}\ .$ (34)
@@ 4. Example I: Probability manifolds from chemical monomolecular triangle reactions
In this section, we present an example of a thermodynamical probability manifold from a system of monomolecular triangle reactions. We note that this is a classic example of Onsager reciprocal relations [21, Section 3]. We present the Levi-Civita connection for this class of probability manifolds.
@@ 4.1. Reciprocal relations
Assume that the reaction obeys a simple mass-action law. The fraction of molecule $A=1$ changes into molecule $B=2$ in a short time $\Delta t>0$ follows
$Q_{12}\Delta t,$
where $Q_{12}$ is a constant reaction rate. It means that the rates of change for the density of three species $A$, $B$, $C$, denoted by $p_{1}$, $p_{2}$, $p_{3}$, satisfy the following master equation system:
$\left\{\begin{aligned} &\frac{dp_{1}}{dt}=Q_{21}p_{2}+Q_{31}p_{3}-(Q_{12}+Q_{13})p_{1}\\ &\frac{dp_{2}}{dt}=Q_{12}p_{1}+Q_{32}p_{3}-(Q_{21}+Q_{23})p_{2}\\ &\frac{dp_{3}}{dt}=Q_{13}p_{1}+Q_{23}p_{2}-(Q_{31}+Q_{32})p_{3}.\end{aligned}\right$. (39)
@@ 4.2. Levi-Civita connection in probability manifolds
$\begin{split}&\langle\bar{\nabla}_{\mathsf{V}_{\Phi_{1}}}\mathsf{V}_{\Phi_{2}},\mathsf{V}_{\Phi_{3}}\rangle\\ =&\frac{\theta_{12}}{2}\Big\{(\nabla_{\omega}\Phi_{1})_{12}(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{3},p))_{12}+(\nabla_{\omega}\Phi_{2})_{12}(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{3},p))_{12}-(\nabla_{\omega}\Phi_{3})_{12}(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{2},p))_{12}\Big\}\\ +&\frac{\theta_{23}}{2}\Big\{(\nabla_{\omega}\Phi_{1})_{23}(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{3},p))_{23}+(\nabla_{\omega}\Phi_{2})_{23}(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{3},p))_{23}-(\nabla_{\omega}\Phi_{3})_{23}(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{2},p))_{23}\Big\}\\ +&\frac{\theta_{13}}{2}\Big\{(\nabla_{\omega}\Phi_{1})_{13}(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{3},p))_{13}+(\nabla_{\omega}\Phi_{2})_{13}(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{3},p))_{13}-(\nabla_{\omega}\Phi_{3})_{13}(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{2},p))_{13}\Big\}.\end{split}$
@@ 5. Example II: Curvatures of probability manifolds on a three-point lattice graph
Again, denote the probability simplex set as $\Delta_{3}$. Given a vector $\Phi\in\mathbb{R}^{3}$ and a point $p\in\Delta_{3}$, consider the metric $\mathbf{g}$ satisfying
@@
We simply notations that $\theta_{1}(p):=\theta_{12}(p)$, and $\theta_{2}(p):=\theta_{23}(p)$. There is a particular coordinate for $\Delta_{3}$, which simplifies the geometric calculations. Denote the cumulative distribution function (CDF) on discrete states, such that
@@
We next present analytical formulas for Riemannian curvatures.
@@
$\bar{R}_{11}(p)=\bar{K}_{12}(p)\theta_{2}(p),\quad\bar{R}_{22}(p)=\bar{K}_{12}(p)\theta_{1}(p),\quad\bar{R}_{12}=0.$
@@
$\left\{\begin{aligned} &\bar{\Gamma}_{11}^{1}=-\frac{1}{2}\partial_{1}\log\theta_{1},\quad\bar{\Gamma}_{11}^{2}=\frac{1}{2}\frac{\theta_{2}}{\theta_{1}}\partial_{2}\log\theta_{1};\\ &\bar{\Gamma}_{22}^{1}=\frac{1}{2}\frac{\theta_{1}}{\theta_{2}}\partial_{1}\log\theta_{2},\quad\bar{\Gamma}^{2}_{22}=-\partial_{2}\log\theta_{2};\\ &\bar{\Gamma}_{12}^{1}=-\frac{1}{2}\partial_{2}\log\theta_{1},\quad\bar{\Gamma}_{12}^{2}=-\frac{1}{2}\partial_{1}\log\theta_{2}.\end{aligned}\right$.
@@
$\left\{\begin{aligned} &\bar{R}_{121}^{2}=\frac{1}{2}\partial_{11}\log\theta_{2}+\frac{1}{4}\partial_{1}\log\theta_{2}\cdot\partial_{1}\log\frac{\theta_{1}}{\theta_{2}}+\frac{\theta_{2}}{\theta_{1}}\Big[\frac{1}{2}\partial_{22}\log\theta_{1}+\frac{1}{4}\partial_{2}\log\theta_{1}\cdot\partial_{2}\log\frac{\theta_{2}}{\theta_{1}}\Big];\\ &\bar{R}_{212}^{1}=\frac{1}{2}\partial_{22}\log\theta_{1}+\frac{1}{4}\partial_{2}\log\theta_{1}\cdot\partial_{2}\log\frac{\theta_{2}}{\theta_{1}}+\frac{\theta_{1}}{\theta_{2}}\Big[\frac{1}{2}\partial_{11}\log\theta_{2}+\frac{1}{4}\partial_{1}\log\theta_{2}\cdot\partial_{1}\log\frac{\theta_{1}}{\theta_{2}}\Big].\end{aligned}\right$.
@@ 6. Discussions
There are insights and applications from geometric calculations in probability manifolds for differential geometry on discrete domains. One example is the generalization of Gamma calculus [3]. Typical examples include the iterative Gamma and Gamma-2 operators. We derive a generalization towards the classical Gamma calculus from the Hessian operator in the thermodynamical probability manifold; see formula (25). This paper also proposes a generalized Gamma calculus in formula (36). We shall investigate properties of Riemannian curvatures, in particular Ricci curvatures, in thermodynamical probability manifolds, which characterize stochastic dynamical behaviors of complex systems [15].
We also mention that there are generalized Onsager response matrices from nonlinear chemical master equations [21]. They are related to general choices of $f$-divergence functions and physical systems in generalized Onsager reciprocal relations. Typical examples include reaction-diffusion equations on discrete domains and dynamical density functional theories on lattices [26]. Systemic computations and spectral analysis from geometric calculations in probability manifolds are essential in understanding dynamical behaviors and fluctuations of thermodynamics. In addition, there are recent studies in non-equilibrium dynamics, from Macroscopic fluctuation theory (MFT) [4] and General Equation for Non-Equilibrium Reversible-Irreversible Coupling (GENERIC) [7]. These studies formulate additional nonreciprocal structures within the Onsager principles, including the conservation of energy and the dissipation of Lyapunov functions [8]. In future works, we leave the geometric calculations and analysis for non-reversible or nonreciprocal relation systems. We shall apply Riemannian curvature in probability manifolds to analyze and design numerical algorithms for physical or chemical models with applications to dynamic Brain health issues.
Acknowledgements. W. Li’s work is supported by the AFOSR YIP award No. FA9550-23-1-0087, NSF RTG: 2038080, and NSF DMS: 2245097.
@@
$\left\{\begin{aligned} &\partial_{1}\log\theta_{1}=\beta(\frac{1}{x_{1}}-\frac{1}{x_{2}-x_{1}}),\quad\partial_{1}\log\theta_{2}=-\frac{\beta}{x_{2}-x_{1}};\\ &\partial_{2}\log\theta_{1}=\frac{\beta}{x_{2}-x_{1}},\quad\partial_{2}\log\theta_{2}=\beta(\frac{1}{x_{2}-x_{1}}-\frac{1}{1-x_{2}});\\ &\partial_{22}\log\theta_{1}=-\frac{\beta}{(x_{2}-x_{1})^{2}},\quad\partial_{11}\log\theta_{2}=-\frac{\beta}{(x_{2}-x_{1})^{2}}.\end{aligned}\right$.

### v4

@@ 3.1. Levi-Civita connection, parallel transport and geodesics
Given $\Phi\in\mathbb{R}^{n}$, we define a vector field $\mathsf{V}_{\Phi}:=L(\theta)\Phi\in{T}_{p}\mathcal{P}_{+}$. Suppose $F\in C^{\infty}(\mathcal{P}_{+};\mathbb{R})$. Denote the directional derivative of $F$ at direction $\mathsf{V}_{\Phi}$ as
@@
$\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}]=L(\mathsf{V}_{\Phi_{1}}\theta)\Phi_{2}-L(V_{\Phi_{2}}\theta)\Phi_{1}.$ (13)
@@
where the last equality is based on the fact that $\nabla^{2}_{pp}F$ is a symmetric matrix. In detail,
@@
$\bar{\nabla}_{\mathsf{V}_{\Phi_{1}}}\mathsf{V}_{\Phi_{2}}=\frac{1}{2}\Big(L(\mathsf{V}_{\Phi_{1}}\theta)\Phi_{2}-L(\mathsf{V}_{\Phi_{2}}\theta)\Phi_{1}+L(\theta)\Gamma(\Phi_{1},\Phi_{2},p)\Big).$ (14)
@@
We next compute the parallel transport. Let $\gamma\colon[0,T]\rightarrow\mathcal{P}_{+}$ be a smooth curve, with a parameter $t>0$. Denote $\mathsf{V}_{\Phi}$ as the tangent direction of the curve $\gamma(t)$ at time $t$. I.e., $\frac{d\gamma(t)}{dt}=L(\theta(\gamma(t)))\Phi(t)=\mathsf{V}_{\Phi(t)}$. Consider a smooth vector field $\mathsf{V}_{\eta}$ given by $\eta(t)\in\mathbb{R}^{n}$. Then equation for $\mathsf{V}_{\eta}$ to be parallel along $\gamma(t)$ satisfies
@@ 3.2. Riemannian curvature tensor
$(abcd):=\Phi_{a}^{\mathsf{T}}L(\mathsf{V}_{\Phi_{b}}\mathsf{V}_{\Phi_{c}}\theta)\Phi_{d}\ .$ (34)
@@ 4. Example I: Probability manifolds from chemical monomolecular triangle reactions
In this section, we present an example of a thermodynamical probability manifold from a system of monomolecular triangle reactions. We note that this is a classic example of Onsager reciprocal relations [22, Section 3]. We present the Levi-Civita connection for this probability manifold.
@@ 4.1. Reciprocal relations
Assume that the reaction obeys a simple mass-action law. The fraction of molecule $A=1$ changes into molecule $B=2$ in a short time $\Delta t>0$ follows $Q_{12}\Delta t$, where $Q_{12}$ is a constant reaction rate. It means that the rates of change for the density of three species $A$, $B$, $C$, denoted by $p_{1}$, $p_{2}$, $p_{3}$, satisfy the following master equation system:
$\left\{\begin{aligned} &\frac{dp_{1}}{dt}=Q_{21}p_{2}+Q_{31}p_{3}-(Q_{12}+Q_{13})p_{1},\\ &\frac{dp_{2}}{dt}=Q_{12}p_{1}+Q_{32}p_{3}-(Q_{21}+Q_{23})p_{2},\\ &\frac{dp_{3}}{dt}=Q_{13}p_{1}+Q_{23}p_{2}-(Q_{31}+Q_{32})p_{3}.\end{aligned}\right$. (39)
@@ 4.2. Levi-Civita connection in probability manifolds
$\begin{split}&\langle\bar{\nabla}_{\mathsf{V}_{\Phi_{1}}}\mathsf{V}_{\Phi_{2}},\mathsf{V}_{\Phi_{3}}\rangle\\ =&\frac{\theta_{12}}{2}\Big\{(\nabla_{\omega}\Phi_{1})_{12}(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{3},p))_{12}+(\nabla_{\omega}\Phi_{2})_{12}(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{3},p))_{12}\\ &\qquad-(\nabla_{\omega}\Phi_{3})_{12}(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{2},p))_{12}\Big\}\\ +&\frac{\theta_{23}}{2}\Big\{(\nabla_{\omega}\Phi_{1})_{23}(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{3},p))_{23}+(\nabla_{\omega}\Phi_{2})_{23}(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{3},p))_{23}\\ &\qquad-(\nabla_{\omega}\Phi_{3})_{23}(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{2},p))_{23}\Big\}\\ +&\frac{\theta_{13}}{2}\Big\{(\nabla_{\omega}\Phi_{1})_{13}(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{3},p))_{13}+(\nabla_{\omega}\Phi_{2})_{13}(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{3},p))_{13}\\ &\qquad-(\nabla_{\omega}\Phi_{3})_{13}(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{2},p))_{13}\Big\}.\end{split}$
@@ 5. Example II: Curvatures of probability manifolds on a three-point lattice graph
Again, denote the probability simplex set as $\Delta_{3}$. We simply notations that $\theta_{1}(p):=\theta_{12}(p)$, and $\theta_{2}(p):=\theta_{23}(p)$. Given a vector $\Phi\in\mathbb{R}^{3}$ and a point $p\in\Delta_{3}$, consider the metric $\mathbf{g}$ satisfying
@@
There is a particular coordinate for $\Delta_{3}$, which simplifies the geometric calculations. Denote the cumulative distribution function (CDF) on discrete states, such that
@@
We next derive formulas for Riemannian curvatures.
@@
$\bar{R}_{11}(p)=\bar{K}_{12}(p)\theta_{2}(p),\quad\bar{R}_{22}(p)=\bar{K}_{12}(p)\theta_{1}(p),\quad\bar{R}_{12}(p)=0.$
@@
$\left\{\begin{aligned} &\bar{\Gamma}_{11}^{1}=-\frac{1}{2}\partial_{1}\log\theta_{1},\quad\bar{\Gamma}_{11}^{2}=\frac{1}{2}\frac{\theta_{2}}{\theta_{1}}\partial_{2}\log\theta_{1},\\ &\bar{\Gamma}_{22}^{1}=\frac{1}{2}\frac{\theta_{1}}{\theta_{2}}\partial_{1}\log\theta_{2},\quad\bar{\Gamma}^{2}_{22}=-\partial_{2}\log\theta_{2},\\ &\bar{\Gamma}_{12}^{1}=-\frac{1}{2}\partial_{2}\log\theta_{1},\quad\bar{\Gamma}_{12}^{2}=-\frac{1}{2}\partial_{1}\log\theta_{2}.\end{aligned}\right$.
@@
$\left\{\begin{aligned} &\bar{R}_{121}^{2}=\frac{1}{2}\partial_{11}\log\theta_{2}+\frac{1}{4}\partial_{1}\log\theta_{2}\cdot\partial_{1}\log\frac{\theta_{1}}{\theta_{2}}+\frac{\theta_{2}}{\theta_{1}}\Big[\frac{1}{2}\partial_{22}\log\theta_{1}+\frac{1}{4}\partial_{2}\log\theta_{1}\cdot\partial_{2}\log\frac{\theta_{2}}{\theta_{1}}\Big],\\ &\bar{R}_{212}^{1}=\frac{1}{2}\partial_{22}\log\theta_{1}+\frac{1}{4}\partial_{2}\log\theta_{1}\cdot\partial_{2}\log\frac{\theta_{2}}{\theta_{1}}+\frac{\theta_{1}}{\theta_{2}}\Big[\frac{1}{2}\partial_{11}\log\theta_{2}+\frac{1}{4}\partial_{1}\log\theta_{2}\cdot\partial_{1}\log\frac{\theta_{1}}{\theta_{2}}\Big].\end{aligned}\right$.
@@ 6. Discussions
There are insights and applications from geometric calculations in probability manifolds for studying “geometry” on discrete domains. One example is the generalization of Gamma calculus [4]. Typical examples include the iterative Gamma and Gamma-2 operators. We derive a generalization towards the classical Gamma calculus from the Hessian operator in the thermodynamical probability manifold; see formula (25). This paper also proposes a generalized Gamma calculus in formula (36). We shall investigate properties of Riemannian curvatures, in particular Ricci curvatures, in probability manifolds, which characterize stochastic dynamical behaviors of complex systems [16].
We also mention that there are generalized Onsager response matrices from nonlinear chemical master equations [22]. They are related to general choices of $f$-divergence functions and physical systems in generalized Onsager reciprocal relations. Typical examples include reaction-diffusion equations on discrete domains and dynamical density functional theories on lattices [27]. Systemic computations and spectral analysis from geometric calculations in probability manifolds are essential in understanding dynamical behaviors and fluctuations of thermodynamics. In addition, there are recent studies in non-equilibrium dynamics, from Macroscopic fluctuation theory (MFT) [5] and General Equation for Non-Equilibrium Reversible-Irreversible Coupling (GENERIC) [8]. These studies formulate additional nonreciprocal structures within the Onsager principles, including the conservation of energy and the dissipation of Lyapunov functions [9]. In future works, we leave the geometric calculations and analysis for non-reversible or nonreciprocal relation systems. We shall apply Riemannian curvature on probability manifolds to analyze and design numerical algorithms for these systems.
Acknowledgements. W. Li’s work was supported by the AFOSR YIP Award (FA9550-23-1-0087), NSF RTG (2038080), NSF DMS (2245097), and the McCausland Faculty Fellowship at the University of South Carolina.
@@
$\left\{\begin{aligned} &\partial_{1}\log\theta_{1}=\beta(\frac{1}{x_{1}}-\frac{1}{x_{2}-x_{1}}),\quad\partial_{1}\log\theta_{2}=-\frac{\beta}{x_{2}-x_{1}},\\ &\partial_{2}\log\theta_{1}=\frac{\beta}{x_{2}-x_{1}},\quad\partial_{2}\log\theta_{2}=\beta(\frac{1}{x_{2}-x_{1}}-\frac{1}{1-x_{2}}),\\ &\partial_{22}\log\theta_{1}=-\frac{\beta}{(x_{2}-x_{1})^{2}},\quad\partial_{11}\log\theta_{2}=-\frac{\beta}{(x_{2}-x_{1})^{2}}.\end{aligned}\right$.